OpenStudy (anonymous):

What is the height, in meters, of the river dam? A.145 cos 65° B. 145 tan 65° C. 145/tan65° D. 145/cos65°

OpenStudy (anonymous):

The picture shows a portion of the river dam.

OpenStudy (hoblos):

tan65 = 145/h h=145/tan65°